# A dimension reduction approach to edge weight estimation for use in spatial models

### v1

@@ 3 Method
Equations 3 and 4 describe two models for areal data, each parameterized by unknown edge weights matrix W. Given a spatial domain representable by graph $G=(V,E)$ having $p$ nodes and $q$ edges, let w be the $q$-length vector containing the unknown weights associated with the edges of $G$. The $p\times p$ edge weights matrix W may be constructed from w and vice versa. We construct the line graph $L(G)$ and obtain the line graph Laplacian $\textbf{L}^{L}$, a $q\times q$ matrix. If $\textbf{V}\mbox{\boldmath$\Lambda$}\textbf{V}^{\top}=\textbf{L}^{L}$ is the eigendecomposition of the LGL, let $\textbf{V}^{k}$ be the $q\times k$ matrix containing the $k$ eigenvectors of $\textbf{L}^{L}$ associated with its $k$ smallest eigenvalues. We can now represent w as follows:
w $=\text{exp}(\textbf{V}^{k}\eta),\;\eta\in\mathbb{R}^{k}$ (6)
@@
Here, $\eta$ is a $k$-length vector of basis coefficients that fully characterize the $p\times p$ edge weights matrix $\textbf{W}(\eta)$. For all $\eta\in\mathbb{R}^{k}$, (and $\sigma^{2}>0$ and $|\kappa|<1$) the CAR covariance matrix $\Sigma=\sigma^{2}(\text{diag}(\textbf{W}(\eta)\textbf{1}_{p})-\kappa\textbf{W}(\eta))^{-1}$ and the GDEF covariance matrix defined in Equation 4 are positive definite matrices. We now discuss some modeling considerations and illustrate properties of this method.
3.1 Contrasting the CAR and GDEF Approaches
We have stated that our framework for parameterizing the edge weights matrix of a given graph may be used when defining both CAR and GDEF models, but provided little discussion so far regarding the implications of using one covariance model or another. CAR models have sparse precision matrices, allowing for computationally efficient likelihood calculations. In contrast, the GDEF model specifies a dense covariance matrix; even when using the proposed basis function parameterization of W, it is virtually impossible to use the GDEF covariance for graphs with over 10,000 nodes due to the impracticality of inverting matrices of that size. The computational advantages of the CAR approach do however come with trade-offs.
Several authors have noted that models which condition only on local or nearest-neighbor structure perform more poorly than those incorporating longer range dependencies (Gramacy and Apley, 2015; Katzfuss and Guinness, 2021; Guinness, 2018). Realizations of spatial random fields produced by CAR models (including ICAR models, which induce perfect correlation between neighboring regions) are considerably less smooth than those generated by the GDEF model, a contrast that is more noticeable on graphs with a large number of nodes. Besag and Mondal (2005) shows that the standard ICAR model on a regular two-dimensional lattice grid converges asymptotically to the distribution of a two-dimensional Brownian motion as grid resolution increases, while Lindgren et al. (2011) shows that the ICAR model may be thought of as a discrete approximation to the limit of a Matérn random field whose range parameter goes to infinity and smoothness goes to zero. Paciorek (2013) demonstrates that CAR models using more than first order adjacency (extending graph structure to second order neighbors and beyond) do not result in meaningfully smoother fields. While computationally efficient, straightforward to implement and widely used, CAR models carry implicit assumptions that may not be desirable in practice.
In contrast to the CAR model, the GDEF model accommodates a wider range of spatial dependence patterns for areal data. Equation 4 defines the GDEF covariance, where $\rho_{\nu}(\cdot)$ is the Matérn correlation function given below:
$\rho_{\nu}(d)=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\sqrt{2\nu}d\right)^{\nu}K_{\nu}\left(\sqrt{2\nu}d\right).$ (7)
Here $d$ is the distance between two observations, $\nu>0$ is a smoothness parameter, often fixed in practice, and $K_{\nu}$ is a modified Bessel function of the second kind of order $\nu$. The exponential covariance function $\sigma^{2}\text{exp}(-d)$ is a special case of the Matérn covariance when $\nu=1/2$; $\nu=3/2$ and $\nu=5/2$ are popular choices for spatial models, and the squared exponential covariance $\sigma^{2}\text{exp}(-d^{2}/2)$ is the limit of Equation 7 as $\nu$ goes to infinity. Those familiar with the Matérn family of covariance functions may note that Equation 7 omits an explicit range parameter. As discussed in Christensen and Hoff (2024), the range parameter of the GDEF model is implicitly defined by the scale of the edge weights matrix, e.g. multiplying W by a factor of two is equivalent to doubling the range parameter, a property we now illustrate.
Figure 4 highlights some distinctions between the CAR and GDEF approaches. Depicted samples were generated over a $30\times 30$ grid from a mean zero normal distribution. Each row corresponds to a different covariance model: the CAR model from Equation 3 with $\kappa=0.95$, the ICAR model with a sum-to-zero constraint (the “smoothest” possible CAR model) and the GDEF model with four different values of $\nu$. In the first column, W was constructed such that $w_{ij}=4$ for all $i\sim j$, with different constants used in subsequent columns. Across all settings the scale parameter $\sigma^{2}$ was set equal to 1. As expected, samples produced by the CAR models and the GDEF model for $\nu=1/2$ are visibly less smooth than those from GDEF models with larger values of $\nu$. The impact that the scale of W has on each model is also notable. As previous stated, the scale of W may thought of as the implicit range parameter within the GDEF; samples with edge weights equal to 4 exhibit slower decay in spatial correlation than those with smaller edge weights. For CAR models, the scale of W doesn’t influence range, but rather the scale of marginal variances. (From Equation 3, it is clear multiplying W by some constant is equivalent to dividing $\sigma^{2}$ by that same constant, indicating that the CAR model as we have defined it is over parameterized, which we address later.)
Figure 4: Samples from $N(\textbf{0},\Sigma)$ over a $30\times 30$ grid with uniform edge weights. Columns correspond to the weight of all nonzero entries in W, while each row corresponds to a different parameterization of $\Sigma$. The first row uses the CAR covariance with $\kappa=0.9$, the second uses the ICAR covariance, and subsequent rows use the GDEF covariance for different values of $\nu$.
The notion of range within CAR models is trickier to characterize than for GDEF models. While the correlation parameter $\kappa$ ostensibly controls the smoothness of the Markov random field and the rate of spatial decay between areal regions, it is highly sensitive to the resolution of the spatial domain. This may be particularly relevant when the CAR model is viewed as the discretization of some continous spatial process. As areal resolution becomes finer, the space of region-averaged continuous spatial processes that could be represented using a CAR or ICAR model shrinks. Figure 5 highlights this, which depicts samples from an ICAR model at different spatial resolutions. If each grid represents the same subset of $\mathbb{R}^{2}$ and samples are viewed as the region-averaged realizations of continuous spatial processes, samples over coarser grids could correspond to smoother underlying processes than those generated at higher resolutions. In general we find that the GDEF model can be used to represent a wider range of spatial dependence patterns, and is less sensitive to how areal units are defined, though there are practical settings in which the computational efficiency of the CAR model may trump these considerations.
Figure 5: Samples from an ICAR model with uniform edge weights at different grid resolutions.
3.2 Impact and interpretation of basis coefficients
@@
The illustrations in Figures 4 and 5 depict samples from distributions defined by uniform edge weights, but a major aim of this article is to show how flexibly defining the edge weights matrix W can lead to more complex and potentially useful covariance structures. As discussed in Christensen and Hoff (2024) regarding the GDEF model, each possible edge weights matrix W corresponds to an embedding of the graph in high-dimensional Euclidean space which is unique up to isometry. The distances between nodes in this embedding may be thought of as the “intrinsic” distances between regions. The GDEF model may be viewed as the graph equivalent of the spatial deformation approach for modeling nonstationary data defined by Sampson and Guttorp (1992), where the deformation function is estimated by placing a prior on the edge weights. This interpretation leads us to say that nodes in regions containing high valued edges are “closer” to each other than regions with low valued edges. While the CAR model does not directly admit a distance based interpretation of W, it can be said of both the CAR and GDEF approaches that edges with large weights correspond to greater connectivity between incident nodes.
@@
Let $\textbf{W}(\eta)$ be the $p\times p$ matrix of edge weights parameterized by $\eta\in\mathbb{R}^{k}$. Equation 6 states that the log of the $q$-length vector containing the unique edge weights in $\textbf{W}(\eta)$ is a linear combination of a collection of $k$ basis functions of length $q$, with $\textbf{V}^{k}$ defined as the eigenvectors of the line graph Laplacian corresponding to the $k$ smallest eigenvectors. We wish to provide some insight into the interpretation of the coefficient vector $\eta=(\eta_{1},...,\eta_{k})^{\prime}$ and illustrate some of the covariance patterns this construction can produce.
@@
As previously noted, the first eigenvector of our basis, $\textbf{v}^{k}_{1}$ is always constant. This means that $\eta_{1}\textbf{v}^{k}_{1}$ is constant and may be viewed as the intercept for the log edge weights. Because $\text{exp}(\textbf{V}^{k}\eta)=\text{exp}(\eta_{1}\textbf{v}^{k}_{1})\odot\text{exp}(\textbf{V}^{k}_{-1}\eta_{-1})$, the vector $\text{exp}(\eta_{1}\textbf{v}^{k}_{1})$ may be viewed as a scaling factor for the edge weights matrix $\textbf{W}(\eta_{-1})$ which is defined using the other $k-1$ basis functions. In the GDEF model, this scaling factor is directly related to the range of spatial correlation, while in the CAR model, the scaling factor is redundant when $\sigma^{2}$ is estimated. As such, we recommend fixing $\eta_{1}=0$ when using the basis function approach to characterizing the edge weights matrix for CAR models. Subsequent coefficients in $\eta$ may be interpreted in context of the eigenvectors with which they are associated. As depicted in Figure 2, the second LGL eigenvector of a $30\times 30$ grid characterizes a contrast between the upper left and lower right corners of the spatial domain. Figure 6 depicts samples from different covariance models defined using only the first two LGL eigenvectors. Across covariance models it can be seen how observations in one corner of the spatial domain are highly correlated, while the effective range of spatial correlation is much smaller in the opposite corner. Spatial covariance patterns such as this do appear in real world settings. Paciorek and Schervish (2006) presents a data set for precipitation in Colorado in which observations east of the Rocky mountains are highly correlated over long distances, whereas spatial correlations between observations in the mountainous western half of the state decay rapidly over short distances.
@@
Figure 6: Samples from $N(\textbf{0},\Sigma)$ over a $30\times 30$ grid using two basis functions with $\eta_{1}=\eta_{2}=50$. Note how spatial correlation decays at different rates across the spatial domain.
@@
Different linear combinations of basis functions can produce a wide array of covariance structures. Figure 7 depicts a linear combination of four eigenvectors that results in covariance where observations at the borders of the spatial domain are highly correlated, while observations in the interior are not. In effect locations at opposite corners of the grid are “closer” together than locations that are both near the center. The more basis functions are included, the more flexible the covariance, though there are generally diminishing returns as the number of eigenvectors used increases due to the higher spatial frequencies exhibited by later eigenvectors. Omitting these high frequency eigenvectors essentially assumes that the edge weights themselves exhibit some degree of spatial smoothness. Intuition on how to select the number of basis functions is provided later in this article.
@@
Figure 7: Illustration of how limited number of basis functions may combine to produce complex covariance structures. $\Sigma$ defined using double exponential covariance.
3.3 Incorporating covariates
We may be interested in the question of whether certain environmental features inhibit or facilitate connectivity between regions in our spatial domain. As such we may wish to model edge weights as a function of environmental covariates. Generally speaking, we will not have covariates that are explictly associated with the edges of our graph. If X is a $p\times r$ matrix containing $r$ features for each of the $p$ nodes in a graph, Hanks and Hooten (2013) suggested defining $\textbf{X}^{E}$, a $q\times r$ feature matrix for the graph’s edges, by averaging the features of adjacent nodes:
$\textbf{x}^{e_{ij}}=\frac{\textbf{x}_{i}+\textbf{x}_{j}}{2}.$ (8)
Here $\textbf{x}^{e_{ij}}$ is the row of $\textbf{X}^{E}$ corresponding to the edge linking nodes $i$ and $j$. The matrix $\textbf{X}^{E}$ could also be constructed treating differences in the features of adjacent regions as a covariate itself. Others have suggested that the length of the shared border between two adjacent regions may be important to consider. However $\textbf{X}^{E}$ is defined, we can augment the basis function model for edge weights in Equation 6 by setting
$\textbf{w}=\text{exp}(\textbf{X}^{E}\psi+\textbf{V}^{k}\eta),\;\psi\in\mathbb{R}^{r},\;\eta\in\mathbb{R}^{k}.$ (9)
[…]
@@ 3.4 Parameter estimation and model fitting
Given graph $G=(V,E)$ with $p$ nodes and $q$ edges, we make repeated observations of the collection of random variables $\{y_{v}:v\in V\}$; each repetition is recorded as $\textbf{y}_{i}\in\mathbb{R}^{p}$ for $i\in 1,...,n$. These observations may be stored in the $n\times p$ data matrix Y, where $y_{ij}$ is the $i$th observation at the $j$th node of $G$. For now, we assume that $\textbf{y}_{i}\overset{iid}{\sim}N_{p}(\textbf{0},\Sigma)$, with $\Sigma$ parameterized by $\theta=(\sigma^{2},\eta)$ for the GDEF model and $\theta=(\sigma^{2},\eta,\kappa)$ for the CAR model. Christensen and Hoff (2024) only considered settings with multiple observations of the spatial process. This was necessary due each edge weight being a unique parameter and the fact $q>p$ in almost all settings. The dimension reduction of the basis function approach to estimation of W allows us to model the covariance of spatial processes with only a single realization ($n=1$).
The maximum likelihood estimator (MLE) for $\theta$ may be obtained using the Newton-Raphson algorithm which requires computation of the first and second derivatives of the log likelihood $\ell(\theta;\textbf{Y})$ with respect to $\theta$. Details on the form of $\nabla_{\ell}(\theta)=\frac{\partial\ell(\theta;\textbf{Y})}{\partial\theta}$ and $\textbf{H}_{\ell}(\theta)=\frac{\partial^{2}\ell(\theta;\textbf{Y})}{\partial\theta^{2}}$ for the models discussed in this article are provided in Appendix 2. Given data Y and an initial set of parameter values $\theta^{0}$ we can maximize the likelihood by iteratively calculating
$\theta^{l}=\theta^{l-1}-\gamma\;\mathbb{E}\left[\textbf{H}_{\ell}(\theta^{l-1})\right]^{-1}\left(\nabla_{\ell}(\theta^{l-1})\right)$ (10)
until $\theta^{l}$ converges to the MLE. Here $\gamma$ is a scaling parameter between 00 and $1$. Smaller $\gamma$ provides more robust optimization, while $\gamma=1$ will generally be faster.
@@
Note when employing the Newton-Raphson algorithm in Equation 10 that we use the expectation of the Hessian matrix $\textbf{H}_{\ell}(\theta)$, which is less computationally intensive to obtain than the Hessian matrix itself. We only need to compute $\textbf{H}_{\ell}(\theta)$ once after obtaining the MLE in order to use the approximation in Equation 11.
We have also used Bayesian estimation techniques to fit the model described in this article. Details on recommend prior choices and MCMC algorithms are also provided in Appendix 2.

### v2

@@ 3 Method
Equations 3 and 4 describe two models for areal data, each parameterized by unknown edge weights matrix W. Let w be the $q$-length vector containing the unknown weights associated with the edges of graph $G$. The $p\times p$ edge weights matrix W may be constructed from w and vice versa. If $\textbf{V}\mbox{\boldmath$\Lambda$}\textbf{V}^{\top}=\textbf{L}^{L}$ is the eigendecomposition of the LGL, let $\textbf{V}^{k}$ be the $q\times k$ matrix containing the $k$ eigenvectors of $\textbf{L}^{L}$ associated with its $k$ smallest eigenvalues. We can now represent w as follows:
w $=\text{exp}(\textbf{V}^{k}\eta),\;\eta\in\mathbb{R}^{k},$ (6)
@@
where $\eta$ is a $k$-length vector of basis coefficients that fully characterizes the $p\times p$ edge weights matrix $\textbf{W}(\eta)$. For all $\eta\in\mathbb{R}^{k}$, (and $\sigma^{2}>0$ and $|\kappa|<1$) the CAR covariance matrix $\Sigma=\sigma^{2}(\text{diag}(\textbf{W}(\eta)\textbf{1}_{p})-\kappa\textbf{W}(\eta))^{-1}$ and the GDEF covariance matrix defined in Equation 4 are positive definite matrices.
A discussion and comparison of the properties of the CAR and GDEF modeling approaches is provided in the supplemental material. In general we find that CAR model is most naturally suited for applications in which the primary focus is the estimation of a dataset’s mean structure; in such settings the CAR framework provides an efficient way to account for spatial random effects that may otherwise obstruct inference on the fixed effects. When inference regarding a dataset’s covariance structure is an important modeling consideration, many authors have noted that the CAR covariance (even when utilizing weighting schemes) has properties that are undesirable or restrictive in many scenarios. See for example Wall, (2004); Besag and Mondal, (2005); Paciorek, (2013). While we find that incorporating the edge weighting scheme of Equation 6 into CAR models improves their flexibility, the primary focus within the remainder of this article will be on the use of the LGL eigenvector basis within the GDEF model.
3.1 Visualizations of method and interpretation of basis coefficients
@@
As discussed in Christensen and Hoff, (2024) regarding the GDEF model, each possible edge weights matrix W corresponds to an embedding of the graph in high-dimensional Euclidean space which is unique up to isometry. The distances between nodes in this embedding may be thought of as the “intrinsic” distances between regions. The GDEF model may be viewed as the graph equivalent of the spatial deformation approach for modeling nonstationary data defined by Sampson and Guttorp, (1992), where the deformation function is obtained by estimating the edge weights. This interpretation leads us to say that nodes in regions containing high valued edges are “closer” to each other than regions with low valued edges. While the CAR model does not directly admit a distance based interpretation of W, it can be said of both the CAR and GDEF approaches that edges with large weights correspond to greater connectivity between incident nodes.
@@
Let $\textbf{W}(\eta)$ be the $p\times p$ matrix of edge weights parameterized by $\eta\in\mathbb{R}^{k}$. Equation 6 states that the log of the $q$-length vector containing the unique edge weights in $\textbf{W}(\eta)$ is a linear combination of a collection of $k$ basis functions of length $q$, with $\textbf{V}^{k}$ defined as the $k$ eigenvectors of the LGL corresponding to its $k$ smallest eigenvalues. We wish to provide some insight into the interpretation of the coefficient vector $\eta=(\eta_{1},...,\eta_{k})^{\prime}$ and illustrate some of the covariance patterns this construction can produce.
@@
Recall that $\eta_{1}\textbf{v}^{k}_{1}$ is constant and may be viewed as the intercept for the log edge weights. Because $\text{exp}(\textbf{V}^{k}\eta)=\text{exp}(\eta_{1}\textbf{v}^{k}_{1})\odot\text{exp}(\textbf{V}^{k}_{-1}\eta_{-1})$, the vector $\text{exp}(\eta_{1}\textbf{v}^{k}_{1})$ may be viewed as a scaling factor for the edge weights matrix $\textbf{W}(\eta_{-1})$ which is defined using the other $k-1$ basis functions. In the GDEF model, this scaling factor is directly related to the range of spatial correlation, while in the CAR model, the scaling factor is redundant when $\sigma^{2}$ is estimated. As such, we recommend fixing $\eta_{1}=0$ when using the basis function approach to characterizing the edge weights matrix for CAR models. Subsequent coefficients in $\eta$ may be interpreted in context of the eigenvectors with which they are associated. As depicted in Figure 2, the second LGL eigenvector of a $\times 30$ 30 grid characterizes a contrast between the top and bottom of the spatial domain. Figure 4 depicts samples from a GDEF covariance model defined using only the first two LGL eigenvectors and smoothness parameter $\nu=3/2$. Across samples it can be seen how observations in the top of the spatial domain are highly correlated, while the effective range of spatial correlation is much smaller on the opposite side. Spatial covariance patterns such as this do appear in real world settings; for example Paciorek and Schervish, (2006) present a data set for precipitation in Colorado in which observations east of the Rocky mountains are highly correlated over long distances, whereas spatial correlations between observations in the mountainous western half of the state decay rapidly over short distances.
@@
Figure 4: Samples from $N(\textbf{0},\Sigma)$ over a $\times 30$ 30 grid using two basis functions with $\eta_{1}=\eta_{2}=50$. Note how spatial correlation decays at different rates across the spatial domain.
@@
Different linear combinations of basis functions can produce a wide array of covariance structures. Figure 5 depicts a linear combination of four eigenvectors that results in covariance where observations at the borders of the spatial domain are highly correlated, while observations in the interior are not. In effect locations at opposite corners of the grid are “closer” together than locations that are both near the center. The more basis functions are included, the more flexible the covariance, though there are generally diminishing returns as the number of eigenvectors used increases due to the higher spatial frequencies exhibited by later eigenvectors. Omitting these high frequency eigenvectors essentially assumes that the edge weights themselves exhibit some degree of spatial smoothness. The question of how to select the number of basis functions is explored later in Section 4, and a discussion regarding the potential inclusion of environmental covariates within our model for edge weights is provided in the supplemental material.
@@
Figure 5: Illustration of how limited number of basis functions may combine to produce complex covariance structures. $\Sigma$ defined using GDEF covariance with $\nu=3/2$.
@@ 3.2 Parameter estimation and model fitting
Much of the spatial deformation literature assumes settings in which there are replicate observations of the spatial process whose covariance is being estimated. Commentary on this assumption is provided in the supplemental material. Should we have replicate observations, we record them as $\textbf{y}_{t}\in\mathbb{R}^{p}$ for $t\in 1,...,n$. These observations may be stored in the $n\times p$ data matrix Y, where $y_{ti}$ is the $t$th observation at the $i$th node of graph $G$. For now, we assume that $\textbf{y}_{t}\overset{iid}{\sim}N_{p}(\textbf{0},\Sigma)$, with $\Sigma$ parameterized by $\theta=(\sigma^{2},\eta)$ for the GDEF model and $\theta=(\sigma^{2},\eta,\kappa)$ for the CAR model. Christensen and Hoff, (2024) only considered settings with multiple observations of the spatial process. This was necessary due to each edge weight being a unique parameter and the fact $q>p$ in almost all settings. Importantly, the dimension reduction of the basis function approach to estimation of W allows us to model the covariance of spatial processes with only a single realization ($n=1$).
The maximum likelihood estimator (MLE) for $\theta$ may be obtained using Fisher scoring which requires computation of the first and second derivatives of the log likelihood $\ell(\theta;\textbf{Y})=-\frac{n}{2}\log|\mbox{\boldmath$\Sigma$}(\mbox{\boldmath$\theta$})|-\frac{1}{2}\sum_{t=1}^{n}\textbf{y}^{T}_{t}\mbox{\boldmath$\Sigma$}^{-1}(\mbox{\boldmath$\theta$})\textbf{y}_{t}$ with respect to $\theta$. Given data Y and an initial set of parameter values $\theta^{0}$ we can maximize the likelihood by iteratively calculating
$\theta^{l}=\theta^{l-1}-\gamma\;\mathbb{E}\left[\textbf{H}_{\ell}(\theta^{l-1})\right]^{-1}\left(\nabla_{\ell}(\theta^{l-1})\right),\hskip 14.22636ptl=1,2,\ldots,$ (7)
until $\theta^{l}$ converges to the MLE. Here $\gamma$ is a scaling parameter between 0 and 111. Smaller $\gamma$ provides more robust optimization, while $\gamma=1$ will generally be faster. Identifying the impact of individual basis function coefficients on the likelihood is fairly challenging, but details on the efficient matrix forms of $\nabla_{\ell}(\theta)=\frac{\partial\ell(\theta;\textbf{Y})}{\partial\theta}$ and $\textbf{H}_{\ell}(\theta)=\frac{\partial^{2}\ell(\theta;\textbf{Y})}{\partial\theta^{2}}$ are provided in the Supplemental material.
@@
Note when employing the algorithm in Equation 7 that we use the expectation of the Hessian matrix $\textbf{H}_{\ell}(\theta)$, which is less computationally intensive to obtain than the Hessian matrix itself. We only need to compute $\textbf{H}_{\ell}(\theta)$ once after obtaining the MLE in order to use the approximation in Equation 8.
3.3 Prediction
The correlation structures of most models for areal data, including the GDEF and especially the CAR models, are sensitive to the structure of the graph on which they are defined (Wall,, 2004; Christensen and Hoff,, 2024). As such, prediction for new or unobserved locations has to be approached more carefully than when using models for point-indexed data to ensure validity.
The log likelihood of $\textbf{y}_{t}$ and its derivatives with respect to $\theta$ are still defined when $\textbf{y}_{t}$ is not fully observed at all nodes within a graph. If one wishes to make predictions for unobserved nodes or new locations, one simply needs to ensure that the graph $G$ used to fit the model is specified such that its nodes and edges are defined for all relevant locations, both where data are observed and where one wishes to predict. Because the LGL basis functions used to characterize the edge weights matrix W are defined over the full network rather than locally, edge structure in unobserved regions of the graph is still estimable. After obtaining estimates for all parameters needed to define $\Sigma(\theta)$, it is straightforward to use the conditional distribution of $\textbf{y}_{\text{predicted}}|\textbf{y}_{\text{observed}}$ to obtain predictions and prediction intervals.